OpenStudy (apollo928):

How would I solve this: At the word “go!” the first person in the chain squeezes the hand of the second person, who in turn immediately squeezes the hand of the third person, and so on. Given the function for time it takes for the signal to transmit around a circle, how long would it take the signal to travel all the way around the world? 91.5 days 65.7 days 102.7 days 83 days Possible answers, help :)

OpenStudy (mathmale):

Distance is involved here. You'd need an approx. measure of the distance around the world. Next, you'd need an approx. measure of how far apart the hands of an average person are. You mention "a function for the time it takes for the signal to transmit around a circle, but you don't provide said function. Seems to me that you'd have to come up with a number of assumptions before you'd be able to start calculating a solution.

OpenStudy (apollo928):

hold on let me type function

OpenStudy (apollo928):

Suppose that the chain could go all the way around the planet. Then the chain’s length would be equal to the circumference of the earth at the equator (about 24,000 miles). Assuming that the average adult arm span is 6 feet, how many people would it take to make this human chain? (Hint: Use 5280 feet = 1 mile.)

jimthompson5910 (jim_thompson5910):

I would start by converting 24000 miles to feet

OpenStudy (apollo928):

I got 21,120,000 for that question but need to know what's the answer for the first question?

jimthompson5910 (jim_thompson5910):

yes you need about 21,120,000 people

jimthompson5910 (jim_thompson5910):

does it says how long it takes for the signal to travel from person to person?

OpenStudy (apollo928):

no

jimthompson5910 (jim_thompson5910):

`Given the function for time it takes for the signal to transmit around a circle` what function are you given?

OpenStudy (apollo928):

There is no given function

jimthompson5910 (jim_thompson5910):

no formulas are written? how strange

OpenStudy (apollo928):

nope

jimthompson5910 (jim_thompson5910):

well it's impossible to answer unless we know how long it takes for one person to receive the signal then to relay it to the next person

OpenStudy (apollo928):

nvm It's s=0.42p

OpenStudy (apollo928):

that would be the function correct?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so you'll plug in p = 21,120,000

OpenStudy (apollo928):

My bad. Didn't realize it at first

jimthompson5910 (jim_thompson5910):

that will give you the total time in seconds. Then you'll convert to days

OpenStudy (apollo928):

how would I convert seconds into days?

jimthompson5910 (jim_thompson5910):

Let's convert from seconds to minutes there are 60 seconds in a minute so \[\Large \left(21,120,000 \ \text{seconds}\right) \times \frac{1 \ \text{minute}}{60 \ \text{seconds}} = \underline{ \ \ \ \ \ \ \ \ } \ \text{minutes}\] fill in the blank. Tell me what you get

OpenStudy (apollo928):

because I got 8,8270,400 when I plugged in 21,120,000 then converted that and got 1021.6481481 for days which doesn't seem right but let me try you're way real quick

OpenStudy (apollo928):

I got 352,000 for min

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

oh wait, I messed up

jimthompson5910 (jim_thompson5910):

I used the wrong value

jimthompson5910 (jim_thompson5910):

Let's convert from seconds to minutes there are 60 seconds in a minute so \[\Large \left(8,870,400 \ \text{seconds}\right) \times \frac{1 \ \text{minute}}{60 \ \text{seconds}} = \underline{ \ \ \ \ \ \ \ \ } \ \text{minutes}\] fill in the blank. Tell me what you get

OpenStudy (apollo928):

I got 147,840 but I don't think It's correct

jimthompson5910 (jim_thompson5910):

yep that's how many minutes it takes \[\Large \left(8,870,400 \ \text{seconds}\right) \times \frac{1 \ \text{minute}}{60 \ \text{seconds}} = \underline{ \ \ \ \ 147,840 \ \ \ \ } \ \text{minutes}\]

jimthompson5910 (jim_thompson5910):

now we must convert from minutes to hours so \[\Large \left(147,840 \ \text{minutes}\right) \times \frac{1 \ \text{hour}}{60 \ \text{minutes}} = \underline{ \ \ \ \ \ \ \ \ } \ \text{hours}\] fill in the blank. Tell me what you get

OpenStudy (apollo928):

2464

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

next convert from hours to days there are 24 hours in a day so \[\Large \left(2,464 \ \text{hours}\right) \times \frac{1 \ \text{day}}{24 \ \text{hours}} = \underline{ \ \ \ \ \ \ \ \ } \ \text{days}\] fill in the blank. Tell me what you get

OpenStudy (apollo928):

So the answer would be 102.6667 which would be 102.7

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (apollo928):

days

OpenStudy (apollo928):

Thank you :)

jimthompson5910 (jim_thompson5910):

no problem